OpenStudy (imamermaid29):

Which functions have a positive rate of change and which have a negative rate of change? Column A 1. 2x − 5y = −15 2. y = −4x + 9 3. x y 2 –10 4 –16 6 –22 4. x y –3 15 –2 11 –1 7 Column B A. positive B. negative

OpenStudy (imamermaid29):

i think the 3rd one is negative and the 4th one is positive

OpenStudy (imamermaid29):

@campbell_st do you know this?

OpenStudy (campbell_st):

yes... its all about the slope of the line. if the slope is positive then the line has positive rate of change. same for negative rate of change

OpenStudy (imamermaid29):

was i right about 3 & 4 ?

OpenStudy (campbell_st):

so for the 1st 2 equations may y the subject of the equation so its in the form y = mx + b m is the slope

OpenStudy (campbell_st):

for 3 and 4 you need the slope formula so just pick to points in 3 if I choose (4, -16) and (6, -22) the slope is \[m = \frac{-22 -(-16)}{6 - 4}\] does that result in a positive number.

OpenStudy (imamermaid29):

yes

OpenStudy (campbell_st):

and in 4 if I pick (-3, 15) and (-1, 7) does \[m = \frac{15 - 7}{-3 -(-1)}\] result in a positive or negative number.

OpenStudy (imamermaid29):

negative

OpenStudy (imamermaid29):

1, negative 2, positive ??